• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 14:41:07 on September 29, 2009

    Entry number 292228

    This entry is a followup to: 291488

    keyword=Right HRS/Compton BCM

    The U3 BCM for the right HRS, which was temporarily changed to U1, has now been restored to U3.

    The FADC Compton BCM, which was T'd off of the U3 signal has been moved. It is now T'd off of the U1 right HRS BCM in the hall, and is still on cable P-5. This signal comes from the counting house patch pannel, port #31.

    A copy of this log entry has been emailed to: mcnulty@jlab.org, rom@jlab.org, kdp2c@virginia.edu, bquinn@cmu.edu